# Deep Transfer Learning: Model Framework and Error Analysis

### v1

@@ 4 Experiment
We substantiate the effectiveness of our method through experimental validation using four image classification datasets and one regression dataset. More details on datasets and hyperparameter setting are shown in Appendix E.
@@
Datasets. The four classification datasets are VLCS [57], PACS [58], TerraIncognita [59] and OfficeHome [60]. All these datasets comprise image data with four distinct domains. The regression dataset is the tool wear dataset from the 2010 PHM Challenge [61], which comprises labeled data with three domains, namely C1, C4, and C6. The original dataset consists of sensing signals sampled at 50 kHz, including cutting force signals, vibration signals in the x, y, and z directions, and acoustic emission signals during processing. We use cutting force signals in three directions as input and the average wear values of the three blades for each tool as output. Since the results for the tool wear were not significant across all domains, we have presented them in Appendix E.
@@
Model selection. In the classification task, we partition data from different domains into training and validation sets in an 8:2 ratio. The transfer process involves the following steps: withholding data from one domain, merging the training sets of the remaining domains to construct the upstream training set and the validation sets to construct the validation set accordingly. For each domain, we vary the hyperparameter $\lambda$, $\tau$ and $\mu$, selecting the model with the highest accuracy on the validation set we construct from the upstream training set. During the downstream task, training is conducted utilizing the training set from the reserved domain, with adjusting the hyperparameter $\kappa,\chi,\zeta$ and $d^{*}$. The model with the highest training accuracy is chosen, and its accuracy on the withheld domain’s validation set is considered as the final accuracy. In the regression task, the distinction lies in partitioning data from various domains into training and validation sets at a 9:1 ratio. Meanwhile, Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Coefficient of determination (R2) [62] are used to evaluate performance of methods.
@@
Table 2: The accuracy of different training methods on typical datasets and their domains
Dataset/Algorithm Accuracy on domain
@@
Comparative methods and their results. We present the test results in Table 2 for the following training methods. ERM-UD (ERM on both upstream and downstream): Training a network without considering task diversity, transferring the feature extractor to the downstream task, and training a linear classifier while freezing the feature extractor. Transfer Invariant Representation (TIR): Adopting and freezing the feature extractor using our upstream training method, followed by training a linear classifier at the downstream task without the $Q$ network. ERM-D (ERM only at downstream): Solely utilizing the pretrained EfficientNet to train the downstream task. Without Independence (WI): Adopting the feature extractor through our upstream training method and incorporating the pretrained EfficientNet while training at the downstream task without imposing any independence constraint between the output of EfficientNet and the transfer featurizer.
Comparison with other works. While our research setting diverges from that of Domain Adaptation, Domain Generalization, and Multi-task Learning as discussed in [66]. And it also differs from single-source domain transfer learning, as discussed in [4]. We conducted comparisons among several Domain Adaptation methods, as outlined in Table 3, utilizing the code from DomainBed [67]. Despite these methods do not impose any constraints on downstream training, the test results demonstrate the superiority of our approach to some extent. Further implementation details regarding the comparison are provided in Appendix E.
Algorithm TerraInc PACS VLCS Office-home Avg
ERM[68] 87.1 $\pm$ 0.1 94.1 $\pm$ 0.3 82.0 $\pm$ 0.6 79.9 $\pm$ 0.3 85.9 $\pm$ 0.3
DRO[69] 87.5 $\pm$ 0.5 94.4 $\pm$ 0.0 82.2 $\pm$ 0.5 80.9 $\pm$ 0.5 85.7 $\pm$ 0.2
IRM[70] 82.2 $\pm$ 0.9 93.6 $\pm$ 0.1 82.0 $\pm$ 0.4 76.2 $\pm$ 0.4 83.5 $\pm$ 0.3
DANN[71] 87.1 $\pm$ 0.2 93.7 $\pm$ 0.4 82.1 $\pm$ 0.9 79.4 $\pm$ 0.1 85.6 $\pm$ 0.3
SagNet[72] 87.3 $\pm$ 0.3 94.0 $\pm$ 0.5 81.7 $\pm$ 0.3 79.5 $\pm$ 0.4 85.7 $\pm$ 0.1
Ours 84.2 $\pm$ 0.2 95.7 $\pm$ 0.4 85.3 $\pm$ 0.1 84.0 $\pm$ 0.3 87.3 $\pm$ 0.1
@@ 5 Conclusion
6 Acknowledgements
Appendices
@@ 6 Acknowledgements
$\mathcal{H}$ to be a function class of features mapping $\mathbb{R}^{d}\to\mathbb{R}^{r}$, $\mathcal{G}$ to be a function class of lipschitz-constrained mapping $\mathbb{R}^{r}\to\mathbb{R}$ and $\mathcal{Q}:\mathbb{R}^{d^{*}}\to\mathbb{R}$ to be a function class during fine-tuning. Also, for any $i\in[p]$, $\mathbb{P}(S=i)>0$, which is a general case and means $n_{i}$ is the same order of n. Without loss of generality, we set the upper bound of random noise as 1: $|\varepsilon_{i}|\leq 1,i\in[2]$.
@@ B.1.1 Proof of error decomposition and approximation decomposition
Here we utilize the metric property of Wasserstein distance and propery of covariance distance in Székely et al. [44, p16,Remark 3]. ∎
@@ B.1.2 Statistical errors
We decompose this part of the error into three components and analyze them separately. The processing for bounding the statistical error of $\mathcal{L}_{1}$ and $\mathcal{L}_{3}$ is almost standard [73, 74]. In the context of analysis restricted to $\mathcal{L}_{2}$, we leveraged the methodological techniques concerning the estimation error of the U-statistic as presented in [75].
@@
The last inequality is due to the Wainwright [76, Exercise 4.7c] and the contraction principle Ledoux and Talagrand [77, Theorem 4.12]. ∎
@@
Condition on $\tilde{D}_{i},i=1,\ldots,n$, let $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$ be the covering number Giné and Nickl [78, page 41] of $\mathcal{H}$ with respect to the empirical distance $d_{D,2}$ at scale of $\delta>0$. Denote $\mathcal{H}_{\delta}$ as the covering set of $\mathcal{H}$ with with cardinality of $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$.
@@
By De la Pena and Giné [79][Corollary 3.2.6], we know that $\xi_{i}$ is all belongs to $\mathcal{L}_{\phi_{2}}$ (a space of Orlicz space with $\phi_{2}(x)=exp(x^{2})$) for their corresponding Orlicz norm [79, page 36] is $C_{0}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$. So it’s still a Sub-Gaussian r.v with parameter $C_{1}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$, and by De la Pena and Giné [79][inequality (4.3.1)] with $\Phi(x)=\exp(\rho x)$ we can get that
@@
where $d_{D,\infty}(\mathbf{h}_{1},\mathbf{h}_{2})=\sum_{i=1}^{n}\frac{1}{n}\|\mathbf{h}_{1}(\mathbf{x}_{i})-\mathbf{h}_{2}(\mathbf{x}_{2})\|_{\infty}$. The proof is base on the proof on Huang et al. [75, Lemma 8.7], and the second inequality is base on the lemma 4.
As for the third last inequality, when estimating the covering number of the family of neural network functions, it is commonly approached by employing concepts such as pseudo dimension or VC dimension. Denote the pseudo dimension of $\mathcal{F}$ by ${\rm Pdim}(\mathcal{F})$, by Anthony et al. [80, Theorem 12.2], for $n\geq{\rm Pdim}(\mathcal{F})$,
@@
Here $B_{\mathcal{F}}$ is the upper bound of funtions in $\mathcal{F}$. However,we can establish its validity regardless of the relative magnitudes of $n$ and ${\rm Pdim}(\mathcal{B}_{pr}(B)\circ\mathcal{H})$, as demonstrated by the proof of Huang et al. [81, Corollary 35].
@@
The first inequality is due to the metric property. The second inequality bounding the term $\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ by Huang et al. [81, lemma 24]. The third inequality is due to the $|\sup_{a}f-\sup_{a}g|\leq\sup_{a}|f-g|$.The second last and last inequality is again by applying chain and symmetrization technique as proof of lemma 3. ∎
@@
Based on [82], the pseudo dimension (VC dimension) of $\mathcal{F}$ with piecewise-linear activation function can be further contained and represented by its depth $L$ and width $W$, i.e., ${\rm Pdim}(\mathcal{F})=O(W^{2}L^{2}\log(W^{2}L)$.
@@ B.2 Proof of Theorem 2
Due to the Theorem 3.3 in [74] and the Theorem 6. For any $\delta>0$, with probability at least $1-\delta$:
@@
C.1 Proof the Theorem 3
proof the Theorem 3.
@@ C.1 Proof the Theorem 3
$+\underbrace{2\mathbb{E}[\sup_{(\mathbf{F},Q_{q,\mathbf{A}})\in\mathcal{B}_{r}(B)\circ\mathcal{Q}_{dd^{*}}}cL(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})-\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})|+\inf_{q\in\mathcal{Q}}\|\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]}_{\mathcal{L}_{5}}.$

### v3

@@ 4 Experiment on four real classification datasets and one regression dataset
We substantiate the effectiveness of our method through experimental validation using four image classification datasets and one regression dataset. The following tables all omit the percentage sign (%) for the sake of simplicity when presenting the classification accuracy. More details on hyperparameter setting are shown in the supplementary material.
Datasets. Among the multi-domain image datasets, PACS (Li et al., 2017), VLCS (Fang et al., 2013), TerraIncognita (Beery et al., 2018) and OfficeHome (Venkateswara et al., 2017) are widely employed. VLCS represents a challenging dataset that is composed of 10,700 images. It encompasses four domains, namely VOC2007 (V), LabelMe (L), Caltech (C), and Sun (S), along with 5 classes. PACS is a dataset comprising 9,900 images, which has 7 classes and 4 domains. TerraIncognita is a geographic dataset consisting of images that cover a wide variety of environments and objects. It consists of 4 domains and 24,700 images. OfficeHome contains images gathered from four distinct domains: Art (A), Clipart (C), Product (P), and Real-world (R), along with corresponding annotations for classification tasks in office environments. It comprises 15,500 images, which are partitioned into 65 object categories. All of these classification datasets are constituted of image data with four distinct domains.
@@
The regression dataset employed is the tool wear dataset from the 2010 PHM Challenge (Li et al., 2009), which is among the rare common datasets utilized for multi-domain regression. This dataset is composed of labeled data spanning three domains, specifically C1, C4, and C6. The original dataset contains sensing signals sampled at a frequency of 50 kHz. These signals encompass cutting force signals, vibration signals in the x, y, and z directions, as well as acoustic emission signals that are generated during the processing stage. We use cutting force signals in three directions as input and the average wear values of the three blades for each tool as output.
@@
Model selection. In the classification task, we partition data from different domains into training and validation sets in an 8:2 ratio. The transfer process involves the following steps: withholding data from one domain, merging the training sets of the remaining domains to construct the upstream training set and the validation sets to construct the validation set accordingly. For each domain, we vary the hyperparameter $\lambda$, $\tau$ and $\mu$, selecting the model with the highest accuracy on the validation set we construct from the upstream training set. During the downstream task, training is conducted utilizing the training set from the reserved domain, with adjusting the hyperparameter $\kappa,\chi,\zeta$ and $d^{*}$. The model with the highest training accuracy is chosen, and its accuracy on the withheld domain’s validation set is considered as the final accuracy. In the regression task, the distinction lies in partitioning data from various domains into training and validation sets at a 9:1 ratio. Meanwhile, Mean Absolute Error (MAE) Root and Mean Square Error (RMSE) are used to evaluate performance of methods.
[…]
Comparison with other works. Two classes were selected from the PACS dataset to compare our method with the KNN-based approach proposed by Cai and Wei (2021), as their method was designed specifically for binary classification. The comparison results are presented in Table 1. The reason why the KNN-based method performs poor is that it is inherently unsuitable for handling high-dimensional data.
Table 1: The accuracy of two selected classes from four domains of PACS
Method/Domain Art Cartoon Photo sketch
KNN-based 62.7 67.5 57.7 49.3
Ours 96.1 97.1 98.7 92.7
Algorithm TerraInc PACS VLCS Office-home Avg
ERM (Vapnik, 1998) 87.1  $\pm$  0.1 94.1  $\pm$  0.3 82.0  $\pm$  0.6 79.9  $\pm$  0.3 85.9  $\pm$  0.3
DRO (Sagawa et al., 2019) 87.5  $\pm$  0.5 94.4  $\pm$  0.0 82.2  $\pm$  0.5 80.9  $\pm$  0.5 85.7  $\pm$  0.2
IRM (Arjovsky et al., 2020) 82.2  $\pm$  0.9 93.6  $\pm$  0.1 82.0  $\pm$  0.4 76.2  $\pm$  0.4 83.5  $\pm$  0.3
DANN (Ganin et al., 2016) 87.1  $\pm$  0.2 93.7  $\pm$  0.4 82.1  $\pm$  0.9 79.4  $\pm$  0.1 85.6  $\pm$  0.3
SagNet (Nam et al., 2021) 87.3  $\pm$  0.3 94.0  $\pm$  0.5 81.7  $\pm$  0.3 79.5  $\pm$  0.4 85.7  $\pm$  0.1
Ours 84.2  $\pm$  0.2 95.7  $\pm$  0.4 85.3  $\pm$  0.1 84.0  $\pm$  0.3 87.3  $\pm$  0.1
While our research setting diverges from that of Domain Adaptation, Domain Generalization, and Multi-task Learning as discussed in Farahani et al. (2021). And it also differs from single-source domain transfer learning, as discussed in Zhuang et al. (2020). To demonstrate the superiority of our learning approach, we conducted comparisons among several renowned Domain Adaptation methods that center around learning effective representations, along with the classical ERM algorithm,, as outlined in Table 2, utilizing the code from DomainBed (Gulrajani and Lopez-Paz, 2020). Despite these methods do not impose any constraints on downstream training, the test results demonstrate the ascendancy of our approach to some extent. Further implementation details regarding the comparison are provided in the supplementary material.
Ablation study and corresponding results. To verify the effectiveness of additional constraints and architecture in our framework. We present the test results in Table LABEL:The_accuracy_of_different_training_methods_on_typical_datasets_and_their_domains(new) for the following training methods. ERM-UD (ERM on both upstream and downstream): Training a network without considering task diversity, transferring the feature extractor to the downstream task, and training a linear classifier while freezing the feature extractor. Transfer Invariant Representation (TIR): Adopting and freezing the feature extractor using our upstream training method, followed by training a linear classifier at the downstream task without the $Q$ network. ERM-D (ERM only at downstream): Solely utilizing the pretrained EfficientNet to train the downstream task. Without Independence (WI): Adopting the feature extractor through our upstream training method and incorporating the pretrained EfficientNet while training at the downstream task without imposing any independence constraint between the output of EfficientNet and the transfer featurizer.
@@
Table 3: The accuracy of ablation study on typical datasets and their domains
Dataset/Alg Accuracy on domain
@@
Results of regression dataset. The results were shown in Table 4. The phrase “C4, C6 -C1” in that represents using the data from domains C4 and C6 as upstream data, and the data from domain C1 as downstream data. The subsequent “C1, C6 -C4” and “C1, C4 -C6” have similar meanings.
Table 4: Fine-tuning accuracy on the Wear dataset
Transfer task/Algorithm C4,C6 -C1 C1,C6 -C4 C1,C4 -C6
Metrics MAE MAE MAE
ERM-UD 8.68 $\pm$ 1.26 5.54 $\pm$ 0.29 2.14 $\pm$ 0.01
Ours 1.22 $\pm$ 0.06 1.47 $\pm$ 0.19 1.23 $\pm$ 0.11
Metrics RMSE RMSE RMSE
ERM-UD 15.10 $\pm$ 0.48 7.36 $\pm$ 0.43 2.77 $\pm$ 0.04
Ours 2.31 $\pm$ 0.15 1.95 $\pm$ 0.21 1.73 $\pm$ 0.12
Table 5: Impact of different representation dimensions on the OfficeHome
@@
A 54.8 $\pm$ 1.7 $57.3\pm 2.8$ 59.4 $\pm$ 1.1 56.9 $\pm$ 0.5 $54.1\pm 1.5$ 50.1 $\pm$ 1.6
C 51.4 $\pm$ 1.4 58.6 $\pm$ 0.9 56.8 $\pm$ 1.4 53.5 $\pm$ 0.5 46.6 $\pm$ 2.3 38.8 $\pm$ 2.3
P 76.1 $\pm$ 1.5 77.9 $\pm$ 1.0 79.4 $\pm$ 2.1 72.5 $\pm$ 1.5 59.7 $\pm$ 1.0 50.3 $\pm$ 0.7
R 71.2 $\pm$ 2.3 77.5 $\pm$ 0.9 76.8 $\pm$ 0.2 74.4 $\pm$ 1.1 65.0 $\pm$ 1.9 59.7 $\pm$ 3.5
Table 6: The Impact of different representation dimensions on the Regression task
[…]
C4,C6 -C1 MAE 6.59 $\pm$ 0.13 4.87 $\pm$ 0.11 $6.20\pm 0.71$ $10.49\pm 0.2$
RMSE 9.17 $\pm$ 0.12 5.89 $\pm$ 0.11 $6.45\pm 0.67$ $15.45\pm 0.20$
C1,C6 -C4 MAE 10.38 $\pm$ 0.04 8.57 $\pm$ 0.17 $10.07\pm 0.38$ $10.33\pm 0.2$
RMSE 15.09 $\pm$ 0.08 13.82 $\pm$ 0.14 $14.98\pm 0.42$ $15.47\pm 0.04$
C1,C4 -C6 MAE 1.64 $\pm$ 0.08 1.41 $\pm$ 0.04 $1.92\pm 0.10$ $2.77\pm 0.23$
RMSE 2.22 $\pm$ 0.04 1.84 $\pm$ 0.01 $2.72\pm 0.13$ $3.45\pm 0.23$
The effect of representation dimension. To determine the optimal dimension of the representation, we conducted experiments on the OfficeHome and the Wear dataset, modifying the dimensionality of the transferred representations based on the TIR training method, which consider the downstream model as (8), with the results shown in Table 5, 6. By comparing the impact of different dimensional representations on downstream tasks, we observed that the accuracy of downstream tasks first increases and then decreases as the dimensionality of the transferred representation increases. This suggests that when the representation dimension is small, it fails to capture enough predictive information, while when the dimension is too large, it introduces unnecessary high dimension information, increasing the cost of estimating a high dimension function during downstream learning. These results align consistently with the result using same technique utilized in our Theorem 3 applied to model (8). Therefore, we selected 1024 and 512 as the dimensionality of the transferred representations for classification tasks and regression tasks respectively .
@@ 5 Conclusion
Learning a pre-trained model with huge data set and then fine-tuning on specific task (Lee et al., 2019; Brown et al., 2020; Hu et al., 2021; Sung et al., 2021; Rücklé et al., 2021; Xu et al., 2021; Liu et al., 2022; Jiang et al., 2022; Zhang et al., 2023; Gao et al., 2023) has empirically proven to be a highly effective strategy in recent years. Our framework for multi-domain transfer learning can also be interpreted along these lines, wherein learning representations from multiple domains can be viewed as pre-training, and refining downstream predictions can be seen as fine-tuning. Consequently, the theoretical results established in this paper also contribute to enhancing our understanding of pre-training and fine-tuning processes.
Supplementary Material to “Deep Transfer Learning: Model Framework and Error Analysis”
Yuling Jiao, Huazhen Lin, Yuchen Luo, Jerry Zhijian Yang
@@ Appendix A Proof of error decomposition and approximation decomposition
$\mathcal{H}$ to be a function class of features mapping $\mathbb{R}^{d}\to\mathbb{R}^{r}$, $\mathcal{G}$ to be a function class of lipschitz-constrained mapping $\mathbb{R}^{r}\to\mathbb{R}$ and $\mathcal{Q}:\mathbb{R}^{d^{*}}\to\mathbb{R}$ to be a function class during fine-tuning.
For any $i\in[p]$, $\mathbb{P}(S=i)>0$, which is a general case and means $n_{i}$ is the same order of n. Without loss of generality, we set the upper bound of random noise as 1: $|\varepsilon_{i}|\leq 1,i\in[2]$.
@@ B.1.1 Proof of error decomposition and approximation decomposition
Here we utilize the metric property of Wasserstein distance and propery of covariance distance in Székely et al. (2007, p16,Remark 3). ∎
@@ B.1.2 Statistical errors
We decompose this part of the error into three components and analyze them separately. The processing for bounding the statistical error of $\mathcal{L}_{1}$ and $\mathcal{L}_{3}$ is almost standard (Anthony and Bartlett, 2009; Mohri et al., 2018). In the context of analysis restricted to $\mathcal{L}_{2}$, we leveraged the methodological techniques concerning the estimation error of the U-statistic as presented in Huang et al. (2024).
@@
The last inequality is due to the Wainwright (2019, Exercise 4.7c) and the contraction principle (Ledoux and Talagrand, 2013, Theorem 4.12). ∎
@@
Condition on $\tilde{D}_{i},i=1,\ldots,n$, let $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$ be the covering number (Giné and Nickl, 2021, page 41) of $\mathcal{H}$ with respect to the empirical distance $d_{D,2}$ at scale of $\delta>0$. Denote $\mathcal{H}_{\delta}$ as the covering set of $\mathcal{H}$ with with cardinality of $\mathfrak{C}(\mathcal{H},d_{D,2},\delta))$.
@@
By De la Pena and Giné (2012)[Corollary 3.2.6], we know that $\xi_{i}$ is all belongs to $\mathcal{L}_{\phi_{2}}$ (a space of Orlicz space with $\phi_{2}(x)=exp(x^{2})$) for their corresponding Orlicz norm (De la Pena and Giné, 2012, page 36) is $C_{0}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$. So it’s still a Sub-Gaussian r.v with parameter $C_{1}\mathrm{E}[\xi_{i}^{2}]^{\frac{1}{2}}$, and by De la Pena and Giné (2012)[inequality (4.3.1)] with $\Phi(x)=\exp(\rho x)$ we can get that
@@
where $d_{D,\infty}(\mathbf{h}_{1},\mathbf{h}_{2})=\sum_{i=1}^{n}\frac{1}{n}\|\mathbf{h}_{1}(\mathbf{x}_{i})-\mathbf{h}_{2}(\mathbf{x}_{2})\|_{\infty}$. The proof is base on the proof on Huang et al. (2024, Lemma 8.7), and the second inequality is base on the lemma 4.
As for the third last inequality, when estimating the covering number of the family of neural network functions, it is commonly approached by employing concepts such as pseudo dimension or VC dimension. Denote the pseudo dimension of $\mathcal{F}$ by ${\rm Pdim}(\mathcal{F})$, by Anthony et al. (1999, Theorem 12.2), for $n\geq{\rm Pdim}(\mathcal{F})$,
@@
Here $B_{\mathcal{F}}$ is the upper bound of funtions in $\mathcal{F}$. However,we can establish its validity regardless of the relative magnitudes of $n$ and ${\rm Pdim}(\mathcal{B}_{pr}(B)\circ\mathcal{H})$, as demonstrated by the proof of Huang et al. (2022, Corollary 35).
@@
The first inequality is due to the metric property. The second inequality bounding the term $\overline{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})-\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ by Huang et al. (2022, lemma 24). The third inequality is due to the $|\sup_{a}f-\sup_{a}g|\leq\sup_{a}|f-g|$.The second last and last inequality is again by applying chain and symmetrization technique as proof of lemma 3. ∎
@@
Based on Bartlett et al. (2019), the pseudo dimension (VC dimension) of $\mathcal{F}$ with piecewise-linear activation function can be further contained and represented by its depth $L$ and width $W$, i.e., ${\rm Pdim}(\mathcal{F})=O(W^{2}L^{2}\log(W^{2}L)$.
@@ B.2 Proof of Theorem 2
Due to the Theorem 3.3 in Mohri et al. (2018) and the Theorem 6. For any $\delta>0$, with probability at least $1-\delta$:
@@
C.1 Proof of the Theorem 3
Proof of the Theorem 3.
@@ C.1 Proof of the Theorem 3
$+\underbrace{2\mathbb{E}[\sup_{(\mathbf{F},Q_{q,\mathbf{A}})\in\mathcal{B}_{r}(B)\circ\mathcal{Q}_{dd^{*}}}\mathcal{L}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})-\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F})|+\inf_{q\in\mathcal{Q}}\|\mathcal{L}(\mathbf{h}^{*},Q_{q,\mathbf{A}^{*}},\mathbf{F}_{T}^{*})-\mathcal{L}(\mathbf{h}^{*},Q^{*},\mathbf{F}_{T}^{*})]}_{\mathcal{L}_{5}}.$